OpenStudy (anonymous):

1/x+1/y=1/z and xy=z what is the average of x and y?

OpenStudy (anonymous):

(x+y)/xy=1/z x+y=1 average = 0.5

OpenStudy (akshay_budhkar):

1/2.....